# Censored extreme value estimation

### v1

@@ Appendix B Proofs of Section 4
$U_{H}(x)=U_{F}\big{(}x^{\alpha_{F}}(\ell_{\tilde{F}}(x))^{-1}\big{)},$ (B.9)
which holds independently of the value of $\gamma_{F}.$ Next, $U_{F}$ is regularly varying with index $\gamma_{F},$ thus by the properties of regularly varying functions, see, e.g., Theorem 1.5.7 (ii) in [8], $U_{H}$ is regularly varying with index $\gamma_{F}\alpha_{F}>0$, which completes the proof in case $\gamma_{F}>0.$
Now assume $\gamma_{F}<0.$ It is easily follows from the second part of Example 2 that if $\alpha_{F}>0$ then $\tau_{F}\leq\tau_{G}$ (to prove this, the condition $\gamma_{G}<0$ is not necessary). Thus, $\tau_{H}=\tau_{F}.$ By (B.9), we derive
@@
The proof relies on the proofs of Theorem 3.1, Theorem 3.3, Theorem 3.4, and Theorem 3.6 under replacement of $\{Z_{i}\}$ with $\{\xi_{i}\},$$\{Z_{n-i,n}/Z_{n-k,n}\}$ with $\{f(\xi_{n-i,n}/\xi_{n-k,n},\xi_{n-k,n})\},$ and $F,$ $G,$ $H$ etc. with their tilde-versions. We also use $g_{1}$ and $g_{2}$ to bound the argument of $\vartheta$ in the analogue of (A.8) and similar situations instead of Potter’s bounds. Consequently, we restrict ourselves to indicating the differences.
First, Lemma A.1 also holds for pairs the $\{(\xi_{i}^{\ast},\delta_{i}^{\ast})\}.$ Indeed, since $H$ is eventually continuous, for large $n$ we get $\xi_{n-i+1,n}=(1-H(Z_{n-i+1,n}))^{-1},$ $i=1,\ldots,k.$ Hence we can choose
$\xi_{i}^{\ast}=\frac{1-H(t)}{1-H(tV_{i}^{\ast})},\quad i=1,\ldots,k,$
where the joint distribution of $\{\xi^{\ast}_{k-i+1,k}\}_{i=1}^{k}$ coincides with the joint distribution of $\{\xi_{n-i+1,n}/\xi_{n-k,n}\}_{i=1}^{k},$ and therefore pairs $\{(\xi_{i}^{\ast},\delta_{i}^{\ast})\}$ are i.i.d as well.
The second difference relates to the argument used in the proof of Theorem 3.1 to show boundedness of certain integrals and first applied to prove (A.8). Indeed, we cannot use that argument as it is, since for that we would need to impose some technical conditions on $f,$ $g_{1}$ and $g_{2}$ like monotonicity and continuity. Let us start with the case of zero $\alpha_{F}$. Consider as an example the version of the relation (A.8) and assume $\alpha_{F}=0$. By the monotone density theorem, see e.g. Theorem 1.7.2 in [8],
@@
$(\tilde{F}^{t})^{\prime}(w)=\frac{\tilde{F}^{\prime}(wt)}{\tilde{F}^{\prime}(t)}\cdot\frac{t\tilde{F}^{\prime}(t)}{1-\tilde{F}(t)}\leq w^{-1+\delta},\quad w\geq 1.$ (B.10)
@@
$|(\tilde{F}^{t})^{\prime}(w)|\leq Aw^{-\alpha_{F}-1+\delta},\quad w\geq 1,$ (B.11)
@@
By (B.10-B.11) and Potter’s bound applied to $1-\tilde{G}^{t},$ we derive for large $t$ and some $C_{1},C_{2},\delta^{\prime}>0$ similarly to (A.9)
@@
$\quad+\,C_{2}\big{|}\bar{\vartheta}(f(w,t))(\tilde{F}^{t})^{\prime}(w)\big{|}\log(w)I(w>1+\delta^{\prime})$
$\quad\leq C_{1}+A\,C_{2}\max\big{(}\bar{\vartheta}(g_{1}(w)),\bar{\vartheta}(g_{2}(w))\big{)}\log(w)\,w^{-\alpha_{F}-1+\delta}.$ (B.12)
The latter function is integrable by Condition 6, and so the dominated convergence theorem can be applied to prove the convergence of $\tilde{\theta}^{t}.$
@@
$\sqrt{k}\big{(}\mu_{f}(\xi_{n-k,n})-\mu_{f}(\infty)\big{)}\stackrel{{\scriptstyle P}}{{\to}}\mu.$ (B.13)
@@
actually coincides with the proof of Theorem 3.9 under replacement of $\varphi,$ $\gamma_{F},$ $F^{\circ},$ etc. with $\vartheta_{g},$ $1/\alpha_{F},$ $\tilde{F}^{\circ},$ etc., respectively. So, let us consider $I_{1}.$ Similarly to (B.7- B.8), we have by (B.11) for large $n$ and some $C,C_{1}>0$,
@@
$\leq$ $C\sqrt{k}\int\max\big{(}\bar{\vartheta}^{\prime}(g_{1}(y)),\bar{\vartheta}^{\prime}(g_{2}(y))\big{)}|f(y,\xi_{n-k,n})-g(y)|\tilde{F}^{\xi_{n-k,n}}(\mathop{}\!\mathrm{d}y)$
$\leq$ $C_{1}\sqrt{k}|\hat{a}(\xi_{n-k,n})|\cdot\int\hat{g}(y)\max\big{(}\bar{\vartheta}^{\prime}(g_{1}(y)),\bar{\vartheta}^{\prime}(g_{2}(y))\big{)}y^{-\alpha_{F}-1+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}y,$
where the last expression is finite by the relation $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\hat{\lambda}$ and Condition 7 and vanishes if $\hat{\lambda}=0,$ which completes the proof for this case. Moreover, (B) implies
$\frac{1}{|\hat{a}(t)|}\Big{|}\big{[}\vartheta(f(y,t))-\vartheta(g(y))\big{]}(\tilde{F}^{t})^{\prime}(y)\Big{|}\leq C_{1}\hat{g}(y)\max\big{(}\bar{\vartheta}^{\prime}(g_{1}(y)),\bar{\vartheta}^{\prime}(g_{2}(y))\big{)}y^{-\alpha_{F}-1+\varepsilon^{\prime}}.$ (B.15)
@@
where $t^{\ast}$ lies between $f(y,t)$ and $g(y).$ The regular variation property of $(\tilde{F})^{\prime}$ and a.e. continuity of $\vartheta^{\prime}$ together with the continuous mapping theorem imply for every $y\geq 1$
$\vartheta^{\prime}(t^{\ast})\stackrel{{\scriptstyle a.e.}}{{\to}}\vartheta^{\prime}(g(y)),\quad(\tilde{F}^{t})^{\prime}(y)\to(\tilde{F}^{\circ})^{\prime}(y),\quad t\to\infty,$
where $\stackrel{{\scriptstyle a.e.}}{{\to}}$ denotes convergence almost everywhere. Therefore, by (B-B.15), the dominated convergence theorem and Condition 4,
@@
Appendix C Proofs of Section 5
@@
First, let us consider the case $\gamma_{F}>0.$ Then $\alpha_{F}$ is positive and we derive by Proposition 4.10 that $G$ and $H$ have regularly varying tails as well, and thus Theorem 3.4 can be applied. The assumption $\alpha_{F}>1/2$ (see also the remark below Theorem 5.1) implies Condition 1 for the functions $\varphi_{\ell}(x)=(\log x)^{\ell},$ $\ell=1,2,$ hence the assumptions of Theorem 3.4 hold for these two functions, obtaining
@@
The proof for the case $\gamma_{F}\leq 0$ is a bit more challenging and is based on the application of Theorem 4.11. Note that since $\alpha_{F}>0$ Proposition 4.10 implies $\gamma_{H}\leq 0$ and $x^{\ast}_{H}=x_{F}^{\ast}.$ As in Section 4.2, let us consider a candidate function
$f(x,t)=\frac{\log U_{H}(tx)-\log U_{H}(t)}{a_{H}(t)/U_{H}(t)},$ (C.1)
and therefore by (4.15) $g(x)=h_{\gamma_{H}}(x),$ where $h_{\gamma}$ is defined by (3.16). The candidates for $g_{1}$ and $g_{2}$ can be found in Section 4.2 as well, see (4.16) for the case $\gamma_{F}<0$ (one should replace $\gamma$ with $\gamma_{H}$ there), whereas for $\gamma_{F}=0$ one can use 0 and $\delta x^{\delta}$ as $g_{1}$ and $g_{2},$ respectively, for some $\delta\in(0,1)$ since $f(x,t)$ is increasing with respect to $x,$ see Remark 4.3.
Consider $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}.$ By the relation $\alpha_{F}+\alpha_{G}=1$ we get $\alpha_{G}<1/2$ and thus, Condition 6 holds for $\vartheta_{1}$ and $\vartheta_{2},$ but one should select a properly small $\delta$ in case $\gamma_{F}=0.$ Next, the condition on existence of eventually monotone density for $\tilde{F}$ can be omitted because it is only needed to prove (B.12) and similar relations. Since $f(w,t)$ is monotone by $w$ and $\tilde{F}$ is regularly varying, the argument similar to the proof of (A.8) can be used to prove the convergence of $\tilde{\theta}^{t}$ and similar expressions. Therefore, applying Theorem 4.11 with necessary changes indicated above, we get
@@
To prove the lemma, we require a two-dimensional analogue of Theorem 3.6, since then all other results of Section 3 concern convergence in probability and hence they are extended to the two-dimensional case straightforwardly. Thus, assume two functions $\varphi^{(1)}$ and $\varphi^{(2)}$ satisfying the assumptions of Theorem 3.6. Conducting the proof of Theorem 3.1 and Theorem 3.3 for the pair $\big{(}S_{k,n}(\varphi^{(1)}),S_{k,n}(\varphi^{(2)})\big{)}$ together, we get similarly to (A.29) the following decomposition:
@@
where for every $j\in\{1,\ldots,k\}$ $W_{j}^{(1)}$ and $W_{j}^{(2)}$ given $Z_{n-k,n}=t$ are the functions of $V^{\ast}_{j}$ and $\delta_{j}^{\ast}$ defined similarly to (A.30) by replacing $V^{\ast},\delta^{\ast}$ with $V^{\ast}_{j},\delta^{\ast}_{j}$ and $\varphi$ with $\varphi^{(1)}$ and $\varphi^{(2)},$ respectively. So $\big{\{}(W_{j}^{(1)},W_{j}^{(2)})\big{\}}_{j=1}^{k}$ are i.i.d. given $Z_{n-k,n}=t,$ and both $\sqrt{k}r^{(1)}_{k,n},$ $\sqrt{k}r^{(2)}_{k,n}$ are $\overline{o}_{P}(1)$ and thus do not affect the asymptotic. Following the proof of Theorem 3.6 in two dimensions (in fact, the only required change is to write $\langle\vec{s},\vec{\varphi}\rangle=s^{(1)}\varphi^{(1)}+s^{(2)}\varphi^{(2)}$ instead of $s\varphi$), we get that the vector
@@
The result of (ii) follows immediately from Theorem 4.12, the analogue of Lemma C.1 for $\gamma_{F}=0$ (the proof technique of the two results is virtually identical), and Cramér’s delta method. The only detail worth clarifying is the calculation of the bias in Theorem 4.12. The function $\tilde{g}$ appearing in the bias term in this theorem is equal to $H_{\gamma_{-},\rho^{\prime}}(\cdot)$ by (4.17), where $\gamma_{-}=\rho^{\prime}=0$ by Lemma B.3.16 in [15]. Thus, $\tilde{g}(x)=\frac{1}{2}\log^{2}x$ by (B.3.8) ibid. The remaining part of the proof is straightforward.
@@
Since under the assumptions of Theorem 5.1, Theorem 4.11 holds for $f$ defined by (C.1) and $g=h_{\gamma_{H}}$, it is enough to show (B.13) in our particular case, see the corresponding steps in the proofs of Theorem 4.11 and Theorem 4.12. Denoting for the moment $s=\xi_{n-k,n},$ we have,
$\mu_{f}(s)-\mu_{f}(\infty)$ $=$ $\int_{1}^{\infty}\vartheta(f(y,s))\tilde{F}^{s}(\mathop{}\!\mathrm{d}y)-\int_{1}^{\infty}\vartheta(g(y))\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y)$ (C.3)
@@
where $\alpha_{F}=\gamma_{H}/\gamma_{F}$ by Example 2 and Proposition 4.10. Denote for brevity $q:=a_{H}/U_{H}$ and recall that $\tau:=\tau_{F}=\tau_{H}$ by Proposition 4.10. By Theorem B.2.18 in [15], $q$ can be chosen equal to $-\gamma_{H}(\tau_{F}-U_{H}).$ Let us change the argument of the integral (C.3) to
@@
and the argument of the integral (C.3) to $x=y^{\gamma_{H}}.$ Notice that these changes are natural since the latter ratio tends to $y^{\gamma_{H}}$ by the properties of distributions belonging to the Weibull max-domain of attraction. Then, we can rewrite
@@
First, consider the asymptotic behaviour of $I_{2}.$ Similar to (A.37), notice that
@@
as in the proof of Theorem 3.9. The remaining part of finding the asymptotic behaviour of $I_{2}$ is a repetition of the steps of that proof (we let the analogue of Condition 2 hold) for $\varphi(x)=\vartheta(\chi(x,0)),$ $u(v,\xi)=1-U_{F^{\ast t}}(\xi),$ and replacement of Condition 3 by the second order condition (5.5). Therefore, we derive
@@
Let us analyse the difference within the squared brackets in the latter integral. By Taylor expansion and the mean value theorem, we get uniformly in $x$
$\vartheta_{H}\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}\right)-\vartheta_{H}(x)$ $=$ $\vartheta_{H}^{\prime}(x^{\ast})\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}-x\right)$ (C.5)
@@
$\sqrt{k}|I_{1}|$$\leq$ $\sqrt{k}\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\left|\frac{x}{1-xt/\tau}-x\right|\mathop{}\!\mathrm{d}(F^{\ast t}(1-x))$ (C.6)
@@
where the latter inequality holds by the argument used for proving (A.8). If
$\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\,x^{1-1/\gamma_{F}+\varepsilon}\mathop{}\!\mathrm{d}x<\infty,$ (C.7)
which holds for $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2},$ then the expression in the right-hand side of (C.6) is finite by the relation $\sqrt{k}(\tau-U_{H}(n/k))\to\hat{\lambda}\tau$ and the argument similar to (A.48), and, moreover, vanishes if $\hat{\lambda}=0,$ which completes the proof of (C.4) for this case.
@@
and assume $\vartheta_{H}^{\prime}$ is a.e. continuous (as is for $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}$). Notice that
@@
for small $t$; in the sequel we use these relations when applying the dominated convergence theorem. Given $\tau-Z_{n-k,n}=t,$ $\sqrt{k}I_{1}$ lies between $\sqrt{k}\amsmathbb{E}\big{[}f_{1}(1-U_{F^{\ast t}}(\xi),t)\big{]}$ and $\sqrt{k}\amsmathbb{E}\big{[}f_{2}(1-U_{F^{\ast t}}(\xi),t)\big{]},$ which both converge to $\hat{\lambda}C_{1}(\gamma_{F})$ as $t\downarrow 0.$ Indeed, by the dominated convergence theorem, (C.5-C.6), convergence $U_{F^{\ast t}}(x)\to U_{F^{\circ}}(x)$ and a.e. continuity of $f_{1}(x,t)$ in the first argument
@@
and the same holds for $f_{2}.$ The above together with $\sqrt{k}(1-U_{H}(n/k)/\tau)\to\hat{\lambda}$ and an argument similar to (A.48) gives us (C.4).
@@
Next, repeating the proof of Lemma C.1 with necessary changes, we get the two-dimensional version of (C.8) with the limit distribution $N(\vec{b},\Sigma),$ where $\vec{b}=(b_{1},b_{2})^{T},$ $\Sigma=\|\Sigma_{ij}\|_{i,j=1}^{2},$ $\Sigma_{11}=\sigma_{1}^{2},$ $\Sigma_{22}=\sigma_{2}^{2},$ and

### v5

@@ Appendix D Proofs of Section 5
$U_{H}(x)=U_{F}\big{(}x^{\alpha_{F}}(\ell_{\tilde{F}}(x))^{-1}\big{)},$ (D.2)
which holds independently of the value of $\gamma_{F}.$ Next, $U_{F}$ is regularly varying with index $\gamma_{F},$ thus by the properties of regularly varying functions; see, e.g., Theorem 1.5.7 (ii) in [7], $U_{H}$ is regularly varying with index $\gamma_{F}\alpha_{F}>0$, which completes the proof in case $\gamma_{F}>0.$
Now assume $\gamma_{F}<0.$ It is easily follows from Proposition 5.1 (i) that if $\alpha_{F}>0$ then $\tau_{F}\leq\tau_{G}$ (to prove this, the condition $\gamma_{G}<0$ is not necessary). Thus, $\tau_{H}=\tau_{F}.$ By (D.2), we derive
@@
The proof relies on the proofs of Theorem B.1, Theorem 3.1, Theorem 3.2, and Theorem 3.4 under replacement of $\{Z_{i}\}$ with $\{\xi_{i}\},$$\{Z_{n-i,n}/Z_{n-k,n}\}$ with $\{f(\xi_{n-i,n}/\xi_{n-k,n},\xi_{n-k,n})\},$ and $F,$ $G,$ $H$ etc. with their tilde-versions. We also use $g_{1}$ and $g_{2}$ to bound the argument of $\vartheta$ in the analogue of (B.11) and similar situations instead of Potter’s bounds. Consequently, we restrict ourselves to indicating the differences.
First, Lemma B.2 also holds for the pairs $\{(\xi_{i}^{\ast},\delta_{i}^{t})\}.$ Indeed, since $H$ is eventually continuous, for large $n$ we get $\xi_{n-i+1,n}=(1-H(Z_{n-i+1,n}))^{-1},$ $i=1,\ldots,k.$ Hence we can choose
$\xi_{i}^{\ast}=\frac{1-H(t)}{1-H(tV_{i}^{t})},\quad i=1,\ldots,k,$
where the joint distribution of $\{\xi^{\ast}_{k-i+1,k}\}_{i=1}^{k}$ coincides with the joint distribution of $\{\xi_{n-i+1,n}/\xi_{n-k,n}\}_{i=1}^{k}$ (and thus does not depend on $t$), and therefore pairs $\{(\xi_{i}^{\ast},\delta_{i}^{t})\}$ are i.i.d as well.
The second difference relates to the argument used in the proof of Theorem B.1 to show boundedness of certain integrals and first applied to prove (B.11). Indeed, we cannot use that argument as it is, since for that we would need to impose some technical conditions on $f,$ $g_{1}$ and $g_{2}$ like monotonicity and continuity. Let us start with the case of zero $\alpha_{F}$. Consider as an example the version of the relation (B.11) and assume $\alpha_{F}=0$. By the monotone density theorem, see e.g. Theorem 1.7.2 in [7],
@@
$(\tilde{F}^{t})^{\prime}(w)=\frac{\tilde{F}^{\prime}(wt)}{\tilde{F}^{\prime}(t)}\cdot\frac{t\tilde{F}^{\prime}(t)}{1-\tilde{F}(t)}\leq w^{-1+\delta},\quad w\geq 1.$ (D.3)
@@
$|(\tilde{F}^{t})^{\prime}(w)|\leq Aw^{-\alpha_{F}-1+\delta},\quad w\geq 1,$ (D.4)
@@
By (D.3-D.4) and Potter’s bound applied to $1-\tilde{G}^{t},$ we derive for large $t$ and some $C_{1},C_{2},\delta^{\prime}>0$ similarly to (B.13)
@@
$\quad+\,C_{2}\big{|}\vartheta(f(w,t))(\tilde{F}^{t})^{\prime}(w)\big{|}\log(w)I(w>1+\delta^{\prime})$
$\quad\leq C_{1}+A\,C_{2}\max\big{(}\phi(g_{1}(w)),\phi(g_{2}(w))\big{)}\log(w)\,w^{-\alpha_{F}-1+\delta},$ (D.5)
where $\phi$ is an envelope of $\vartheta$ satisfying Condition 5. The function in the right-hand side of (D.5) is integrable by Condition 5, and so the dominated convergence theorem can be applied to prove the convergence of $\tilde{\theta}^{t}.$
@@
$\sqrt{k}\big{(}\mu_{f}(\xi_{n-k,n})-\mu_{f}(\infty)\big{)}\stackrel{{\scriptstyle P}}{{\to}}\mu.$ (D.6)
@@
actually coincides with the proof of Theorem 3.6 under replacement of $\varphi,$ $\gamma_{F},$ $F^{\circ},$ etc. with $\vartheta_{g},$ $1/\alpha_{F},$ $\tilde{F}^{\circ},$ etc., respectively. Here we also use an analogue of (B.44), for handling the behavior of $\vartheta_{g}^{\prime}$:
$\int_{1}^{\infty}|\vartheta_{g}^{\prime}(x)|\,x^{-\alpha_{F}+\varepsilon}\mathop{}\!\mathrm{d}x<\infty,$
which follows from Condition 5, integration by parts, and continuity and eventual monotonicity of $\vartheta_{g}^{\prime}.$
So, let us consider $I_{1}.$ As in the proofs above, let us take an envelope $\phi$ of $\vartheta^{\prime}$ satisfying Condition 6. Similarly to (C.7- C.8), we have by (D.4) for large $n$ and some $C,C_{1}>0$,
@@
$\leq$ $C\sqrt{k}\int\max\big{(}\phi(g_{1}(y)),\phi(g_{2}(y))\big{)}|f(y,\xi_{n-k,n})-g(y)|\tilde{F}^{\xi_{n-k,n}}(\mathop{}\!\mathrm{d}y)$
$\leq$ $C_{1}\sqrt{k}|\hat{a}(\xi_{n-k,n})|\cdot\int\hat{g}(y)\max\big{(}\phi(g_{1}(y)),\phi(g_{2}(y))\big{)}y^{-\alpha_{F}-1+\varepsilon^{\prime}}\mathop{}\!\mathrm{d}y,$
where the last expression is finite by the relation $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\hat{\lambda}$ and Condition 6 and vanishes if $\hat{\lambda}=0,$ which completes the proof for this case. Moreover, (D) implies
$\frac{1}{|\hat{a}(t)|}\Big{|}\big{[}\vartheta(f(y,t))-\vartheta(g(y))\big{]}(\tilde{F}^{t})^{\prime}(y)\Big{|}$ (D.8)
$\leq C_{1}\hat{g}(y)\max\big{(}\phi(g_{1}(y)),\phi(g_{2}(y))\big{)}y^{-\alpha_{F}-1+\varepsilon^{\prime}}.$
@@
where $t^{\ast}$ lies between $f(y,t)$ and $g(y).$ The regular variation property of $(\tilde{F})^{\prime}$ and continuity of $\vartheta^{\prime}$ together with the continuous mapping theorem imply for every $y\geq 1$
$\vartheta^{\prime}(t^{\ast}){\to}\vartheta^{\prime}(g(y)),\quad(\tilde{F}^{t})^{\prime}(y)\to(\tilde{F}^{\circ})^{\prime}(y),\quad t\to\infty.$
Therefore, by (D-D.8), the dominated convergence theorem and Condition 3,
@@
Appendix E Proofs of Section 6
@@
First, let us consider the case $\gamma_{F}>0.$ Then $\alpha_{F}$ is positive and we derive by Proposition 5.1 (ii) that $G$ and $H$ have regularly varying tails as well, and thus Theorem 3.2 can be applied. The assumption $\alpha_{F}>1/2$ (see also the remark below Theorem 6.1) implies Condition 1 for the functions $\varphi_{\ell}(x)=(\log x)^{\ell},$ $\ell=1,2,$ hence the assumptions of Theorem 3.2 hold for these two functions, obtaining
@@
The proof for the case $\gamma_{F}\leq 0$ is a bit more challenging and is based on the application of Theorem 5.2. Note that since $\alpha_{F}>0$ Proposition 5.1 (ii) implies $\gamma_{H}\leq 0$ and $\tau_{H}=\tau_{F}.$ As in Section A, let us consider a candidate function
$f(x,t)=\frac{\log U_{H}(tx)-\log U_{H}(t)}{a_{H}(t)/U_{H}(t)},$ (E.1)
and therefore by (A.4) $g(x)=h_{\gamma_{H}}(x),$ where $h_{\gamma}$ is defined by (3.11). The candidates for $g_{1}$ and $g_{2}$ can be found in Section A as well; see (A.5) for the case $\gamma_{F}<0$ (one should replace $\gamma$ with $\gamma_{H}$ there), whereas for $\gamma_{F}=0$ one can use 0 and $\delta x^{\delta}$ as $g_{1}$ and $g_{2},$ respectively, for some $\delta\in(0,1)$ since $f(x,t)$ is increasing with respect to $x$.
Consider $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}.$ By the relation $\alpha_{F}+\alpha_{G}=1$ we get $\alpha_{G}<1/2$ and thus, Condition 5 holds for $\vartheta_{1}$ and $\vartheta_{2},$ but one should select a properly small $\delta$ in case $\gamma_{F}=0.$ Next, the condition on existence of eventually monotone density for $\tilde{F}$ can be omitted because it is only needed to prove (D.5) and similar relations. Since $f(w,t)$ is monotone by $w$ and $\tilde{F}$ is regularly varying, the argument similar to the proof of (B.11) can be used to prove the convergence of $\tilde{\theta}^{t}$ and similar expressions. Therefore, applying Theorem 5.2 with the necessary changes indicated above, we get
@@
To prove the lemma, we require a two-dimensional analogue of Theorem 3.4, since then all other results of Section 3 concern convergence in probability and hence they are extended to the two-dimensional case straightforwardly. Thus, assume two functions $\varphi^{(1)}$ and $\varphi^{(2)}$ satisfying the assumptions of Theorem 3.4. Conducting the proof of Theorem B.1 and Theorem 3.1 for the pair $\big{(}S_{k,n}(\varphi^{(1)}),S_{k,n}(\varphi^{(2)})\big{)}$ together, we get similarly to (B.35) the following decomposition:
@@
where for every $j\in\{1,\ldots,k\}$ $W_{j}^{(1)}$ and $W_{j}^{(2)}$ given $Z_{n-k,n}=t$ are the functions of $V^{t}_{j}$ and $\delta_{j}^{t}$ defined similarly to (B.36) by replacing $V^{t},\delta^{t}$ with $V^{t}_{j},\delta^{t}_{j}$ and $\varphi$ with $\varphi^{(1)}$ and $\varphi^{(2)},$ respectively. So $\big{\{}(W_{j}^{(1)},W_{j}^{(2)})\big{\}}_{j=1}^{k}$ are i.i.d. given $Z_{n-k,n}=t,$ and both $\sqrt{k}r^{(1)}_{k,n},$ $\sqrt{k}r^{(2)}_{k,n}$ do not affect the asymptotic by the first argument in the proof of Theorem 3.2. Following the proof of Theorem 3.4 in two dimensions (in fact, the only required change is to write $\langle\vec{s},\vec{\varphi}\rangle=s^{(1)}\varphi^{(1)}+s^{(2)}\varphi^{(2)}$ instead of $s\varphi$), we get that the vector
@@
Note that it is possible to prove assertion (i) by imposing the second-order condition (A.6) instead of Condition 2, as it was done, e.g., in Theorem 3.5.4, [11] in the uncensored case. We avoid such a route to highlight that for $\gamma_{F}>0$ the proof follows quite simply from Theorem 3.6.
The result of (ii) follows immediately from Theorem 5.3, the analogue of Lemma E.1 for $\gamma_{F}=0$ (the proof technique of the two results is virtually identical), and Cramér’s delta method. The only detail worth clarifying is the calculation of the bias in Theorem 5.3. The function $\tilde{g}$ appearing in the bias term in this theorem is equal to $H_{\gamma_{-},\rho^{\prime}}(\cdot)$ by (4.13), where $\gamma_{-}=\rho^{\prime}=0$ by Lemma B.3.16 in [11]. Thus, $\tilde{g}(x)=\frac{1}{2}\log^{2}x$ by (B.3.8) ibid. The remaining part of the proof is straightforward.
@@
Since under the assumptions of Theorem 6.1, Theorem 5.2 holds for $f$ defined by (E.1) and $g=h_{\gamma_{H}}$, it is enough to show (D.6) in our particular case; see the corresponding steps in the proofs of Theorem 5.2 and Theorem 5.3. Denoting for the moment $s=\xi_{n-k,n},$ we have,
$\mu_{f}(s)-\mu_{f}(\infty)$ $=$ $\int_{1}^{\infty}\vartheta(f(y,s))\tilde{F}^{s}(\mathop{}\!\mathrm{d}y)-\int_{1}^{\infty}\vartheta(g(y))\tilde{F}^{\circ}(\mathop{}\!\mathrm{d}y)$ (E.4)
@@
where $\alpha_{F}=\gamma_{H}/\gamma_{F}$ by Proposition 5.1. Denote for brevity $q:=a_{H}/U_{H}$ and recall that $\tau:=\tau_{F}=\tau_{H}$ by Proposition 5.1 (ii). By Theorem B.2.18 in [11], $q$ can be chosen equal to $-\gamma_{H}(\tau_{F}-U_{H}).$ Let us change the argument of the integral (E.4) to
@@
and the argument of the integral (E.4) to $x=y^{\gamma_{H}}.$ Notice that these changes are natural since the latter ratio tends to $y^{\gamma_{H}}$ by the properties of distributions belonging to the Weibull max-domain of attraction. Then, we can rewrite
@@
First, consider the asymptotic behavior of $I_{2}.$ Similar to (B.47), notice that
@@
as in the proof of Theorem 3.6. The remaining part of finding the asymptotic behavior of $I_{2}$ is a repetition of the steps of that proof (we let the analogue of Condition B.44 hold) for $\varphi(x)=\vartheta(\chi(x,0)),$ $u(v,\xi)=1-U_{F^{\ast t}}(\xi),$ and replacement of Condition 2 by the second-order condition (6.5). Therefore, recalling that $t=\tau-Z_{n-k,n},$ we derive
@@
Let us analyze the difference within the squared brackets in the latter integral. By Taylor expansion and the mean value theorem, we get uniformly in $x$
$\vartheta_{H}\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}\right)-\vartheta_{H}(x)$ $=$ $\vartheta_{H}^{\prime}(x^{\ast})\left(\frac{\log(1-xt/\tau)}{\log(1-t/\tau)}-x\right)$ (E.6)
@@
$\sqrt{k}|I_{1}|$$\leq$ $\sqrt{k}\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\left|\frac{x}{1-xt/\tau}-x\right|\mathop{}\!\mathrm{d}(F^{\ast t}(1-x))$ (E.7)
@@
where the latter inequality holds by the argument used for proving (B.11). If
$\int_{0}^{1}|\vartheta_{H}^{\prime}(x)|\,x^{1-1/\gamma_{F}+\varepsilon}\mathop{}\!\mathrm{d}x<\infty,$ (E.8)
which holds for $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2},$ then the expression in the right-hand side of (E.7) is finite by the relation $\sqrt{k}(\tau-U_{H}(n/k))\to\hat{\lambda}\tau$ and the argument similar to (B.58), and, moreover, vanishes if $\hat{\lambda}=0,$ which completes the proof of (E.5) for this case.
@@
and assume $\vartheta_{H}^{\prime}$ is continuous (as is for $\vartheta_{1}(x)=x$ and $\vartheta_{2}(x)=x^{2}$). Notice that
@@
for small $t$; in the sequel we use these relations when applying the dominated convergence theorem. Given $\tau-Z_{n-k,n}=t$ (here we think about $t$ as a real number), $\sqrt{k}I_{1}$ lies between $\sqrt{k}\amsmathbb{E}\big{[}f_{1}(1-U_{F^{\ast t}}(\xi),t)\big{]}$ and $\sqrt{k}\amsmathbb{E}\big{[}f_{2}(1-U_{F^{\ast t}}(\xi),t)\big{]},$ which both converge to $\hat{\lambda}C_{1}(\gamma_{F})$ as $t\downarrow 0.$ Indeed, by the dominated convergence theorem, (E.6-E.7), convergence $U_{F^{\ast t}}(x)\to U_{F^{\circ}}(x)$ and continuity of $f_{1}(x,t)$ in the first argument
@@
and the same holds for $f_{2}.$ The above together with $\sqrt{k}(1-U_{H}(n/k)/\tau)\to\hat{\lambda}$ and an argument similar to (B.58) gives us (E.5).
@@
Next, repeating the proof of Lemma E.1 with necessary changes, we get the two-dimensional version of (E.9) with the limit distribution $N(\vec{b},\Sigma),$ where $\vec{b}=(b_{1},b_{2})^{T},$ $\Sigma=\|\Sigma_{ij}\|_{i,j=1}^{2},$ $\Sigma_{11}=\sigma_{1}^{2},$ $\Sigma_{22}=\sigma_{2}^{2},$ and